If the distance between the point (8,7) and (3,y) is 13units.Find y
How can I find why if they gave me any numbers in order to find y?
Please show me how to solve this problem so if I run into it again I know how to do it

Answer by checkley77(12844)

D=SQRT[(Y2-Y1)^2+(X2-X1)^2]]
13=SQRT[(Y-7)^2+(3-8)^2]
13=SQRT[Y^2-14Y+49-5^2]
13=SQRT[Y^2-14Y+49+25]
13=SQRT[Y^2-14Y+74]
13^2=Y^2-14Y+74
169=Y^2-14Y+74
Y^2-14Y+74-169=0
Y^2-14Y-95
(Y-19)(Y+5)=0
Y-19=0
Y=19 ANS.
Y+5=0
Y=-5 ANS.